### v2

# Position: Simulating Society Requires Simulating Thought

Abstract
Simulating society with large language models (LLMs), we argue, requires more than generating plausible behavior; it demands cognitively grounded reasoning that is structured, revisable, and traceable. LLM-based agents are increasingly used to emulate individual and group behavior, primarily through prompting and supervised fine-tuning. Yet they often lack internal coherence, causal reasoning, and belief traceability, making them unreliable for simulating how people reason, deliberate, and respond to interventions.
To address this, we present a conceptual modeling paradigm, Generative Minds (GenMinds), which draws from cognitive science to support structured belief representations in generative agents. To evaluate such agents, we introduce the RECAP (REconstructing CAusal Paths) framework, a benchmark designed to assess reasoning fidelity via causal traceability, demographic grounding, and intervention consistency. These contributions advance a broader shift: from surface-level mimicry to generative agents that simulate thought—not just language—for social simulations.
*Equal contribution.
Now at Google.
1 Introduction
Over the past two years, LLMs have increasingly become dominant tools for simulating human behavior across language [1, 2], vision [3] and decision-making domains [4].In the field of social simulation, LLMs are now commonly used to emulate public opinions, stakeholder interactions, and policy responses under diverse scenarios [5, 6].
Despite these advances, existing models remain far from cognitively intelligent. They often exhibit shallow reasoning, frequent hallucinations, and limited understanding of commonsense causal relationships in socially-relevant topics such as upzoning, surveillance, or healthcare access [7, 8].
In addition, these models are difficult to adapt to new domains. Fine-tuning LLMs for context-specific simulation requires substantial computational resources and access to large, high-quality datasets, both of which are often scarce in real-world policy settings. Yet effective social simulation demands precisely the opposite: the ability to represent dynamically evolving stakeholder reasoning, grounded in timely and localized information [9, 10].
Moreover, current LLM-based agents remain largely opaque in their internal reasoning under contextual change. It is difficult to determine how much a given policy change or contextual shift influences the agent’s opinions or decisions—let alone why. Most existing systems operate as black boxes, lacking interpretable reasoning traces that reveal how conclusions are formed. This is a critical limitation in the social simulation domain, where understanding how individuals arrive at particular judgments is often more important than the judgments themselves.
While post-hoc analysis of outputs is common, it is insufficient for modeling complex, multi-stakeholder reasoning. Without transparent internal structures, agents cannot support diagnostic explanation, causal attribution, or meaningful intervention—all of which are essential for interpreting public responses to policy and for building trustworthy simulations.
These limitations call for a shift in how we conceptualize generative social simulation—not as behavior mimicry, but as cognitive modeling. This paper takes up that call. Specifically, we propose leveraging ideas from Theory of Mind (ToM) and cognitive science to extract and simulate structured reasoning templates—what we term reasoning traces—rather than simply mimicking human tone or persona [11, 12].
Unlike prompt-driven persona or character approaches that generate “average” group behaviors [13], cognitive models allow agents to represent beliefs, values, and causal assumptions in a compositional manner. This makes it possible to generalize to unseen scenarios, so long as the individual components of the reasoning trace are known. For example, if a stakeholder has previously reasoned about “density” and “transit,” then when asked about a novel “transit-oriented development” policy, the agent can reuse those motifs to simulate beliefs without re-training.
Such modularity is the cornerstone of human cognition, where reasoning is built from reusable, revisable fragments across contexts [14, 15]. It also makes simulation more faithful: interactions between agents, or between agents and environment dynamics, can be represented as separable, inspectable modules—enabling structured simulation at both micro and macro levels [16]. Moreover, by reusing modular reasoning components rather than re-generating full-context reasoning each time, this approach significantly reduces the token budget required during simulation and inference—making it both more interpretable and more computationally efficient.
In this paper, we advocate moving beyond output-level alignment toward aligning the internal reasoning traces of generative agents. Capturing the causal, compositional, and revisable structure of belief formation—what we call reasoning fidelity—is essential to building cognitively faithful agents that simulate not only what people say, but how they think.
To support this argument, we:
  • Theorize reasoning fidelity as a structural alignment problem grounded in cognitive science;
  • Introduce a symbolic-neural framework for simulating belief formation through modular reasoning motifs and causal graphs;
  • Present a methodology for extracting and simulating belief structures from natural language, enabling interpretability, counterfactual reasoning, and domain transfer;
  • Illustrate how current approaches fall short by producing outputs that appear coherent but lack internal consistency, adaptability, or traceability.
In summary, this position paper argues that simulating human society requires more than generating plausible conversations. It requires simulating the structure of human reasoning. By grounding agents in modular belief representations and evaluating them on reasoning fidelity, we take a critical step toward building generative minds, not just generative outputs.
2 Social Simulation: Opportunities and Gaps
Social simulation has emerged as a high-impact use case for LLMs, particularly in policy modeling, urban planning, and behavior forecasting. Traditionally, social science relies on surveys, experiments, and fieldwork to understand individual and group behavior. While effective, these methods are expensive, hard to scale, and often face ethical and logistical challenges. The rise of LLM-driven agents offers a promising alternative—capable of simulating human responses across a wide range of scenarios, roles, and interventions.
Recent studies demonstrate that LLMs can emulate key aspects of human reasoning and decision-making [17, 18, 19, 20], enabling agents that perceive their environment, make context-sensitive decisions, and articulate motivations. When equipped with role-based prompting or persona conditioning [21, 22], these agents exhibit a property known as algorithmic fidelity—the ability to simulate how specific individuals or subgroups might respond in a given situation [23, 24].
Beyond single-agent settings, multi-agent simulations further extend this potential by modeling the interactions, conflicts, and consensus dynamics among synthetic populations. However, faithfully simulating social processes introduces several critical requirements:
  1. 1.
    Fidelity: Simulated agents must remain sensitive to environmental and policy contexts, updating their beliefs and responses accordingly [25, 26].
  2. 2.
    Diversity: Outputs should reflect the heterogeneity of real populations—avoiding “flattening” effects where different demographics collapse into a single voice [25, 27].
  3. 3.
    Scalability: Simulations must scale to the population level to support distributional analysis of public opinion and policy impact [28, 29].
  4. 4.
    Generalization: Models should extrapolate to unseen cases by reusing previously learned reasoning motifs across domains [26, 27].
  5. 5.
    Consistency: A simulated agent should maintain a coherent reasoning trace across different scenarios, with beliefs that evolve predictably rather than being reset at each prompt [29].
While many recent works aim to create more “human-like” agents, few clearly define what this term means—or how such fidelity is to be evaluated. Most existing approaches focus on behavior-level plausibility, often relying on post-hoc justification rather than simulating the underlying structure of belief formation and revision.
In this work, we take a complementary approach by grounding agent simulation in cognitive theory. We argue that modeling reasoning traces—explicit, modular structures of how beliefs are formed and revised—enables more faithful, transparent, and generalizable social simulations. Our aim is not only to guide simulator design but also to provide clearer standards for evaluating claims of human-likeness in generative agents. While our primary focus is social simulation, we believe our approach contributes to a broader shift toward cognitively grounded agent modeling in human-AI systems.
3 Problem Statement: Behavioral Plausibility is Not Enough
Most existing efforts to align LLM-based agents focus on outputs [30]: Do they express coherent stances, mimic preferences, or engage in natural-sounding conversations? This behavior-centric view is reinforced by popular techniques such as reinforcement learning from human feedback (RLHF) [31, 32], persona prompting [33], and chain-of-thought (CoT) generation [17, 34]. These methods optimize for plausibility—not for structural correctness of thought.
This output-centric view overlooks a critical problem: output plausibility is not equivalent to cognitive alignment. In this section, we argue that behavioral fluency is a poor substitute for reasoning fidelity. Without internal representations of belief, mechanisms for revision, or sensitivity to positional diversity, generative agents risk producing superficially plausible outputs that are epistemically hollow. We identify two core challenges—diversity, or the agent’s ability to model distributed and positional human reasoning, and fidelity, or the agent’s capacity for coherent, revisable, and causally grounded belief formation—and show how both are undermined by current architectures and deficient evaluation metrics.
3.1 Diversity: Social Simulation Without Structured Reasoning Fails Downstream
When deployed in real-world contexts—such as civic simulations [35, 36], participatory policy design [37], or stakeholder modeling [38]—generative agents are increasingly tasked with simulating belief dynamics, forecasting societal responses, or engaging in normative deliberation [39, 40, 23, 41, 42, 43]. In these settings, agents that lack internal reasoning fidelity may produce outputs that appear thoughtful, yet encode no coherent decision process beneath the text. This disconnect introduces a set of critical downstream failures:
Flattened Outputs within Demographic Groups.
Identity flattening occurs when agents simulate demographic groups without attending to the intersectional nature of lived experience. Current LLMs (and thus model-based agents) are not trained to reason about how multiple dimensions of identity interact to shape values, risk perceptions, or decision preferences but default to essentialize stereotypical portrayals, producing responses that erase internal diversity within groups, thus systematically misrepresenting marginalized communities and flatten intra-group variation, which often reflects dominant-culture priors embedded in their pretraining data [44, 45]. This leads to epistemic harm: the rich, positional knowledge of real-world stakeholders is replaced with monolithic, decontextualized simulations.
Motivating Example: As Wang et al. document, identity essentialization is common in current LLMs when generative agents are prompted with demographic labels and stereotypical portrayals [44]:
“Examples of identity essentialization from GPT-4, when prompted with the identity of Black woman, include the outputs ‘Hey girl!’, ‘Hey sis,’ and ‘Oh, honey’; compared to White man with ‘Hey buddy,’ ‘Hey, friend!’ and ‘Hey mate.’ Llama-2 for Black women starts most responses with ‘Oh, girl,’ and uses phrases like ‘I’m like, YAASSSSS’ and ‘That’s cray, hunty!’”
Alternative View: Generalization over identity categories is necessary for tractable simulation.
One might argue that abstractions over demographic identities are unavoidable and, in fact, desirable when building simulators at scale considering model tractability [46]. Identity flattening may be viewed as a form of necessary regularization.
Response. Our critique is not about the use of abstraction per se, but the fact that LLMs abstract without modeling the joint distribution of beliefs, values, and positionality conditioned on intersecting variables (e.g., age × race × class × institutional exposure) and how abstraction, subsequently, is operationalized without epistemic grounding [44]. In social simulations, such abstractions introduce bias, undermine group heterogeneity, and lack epistemic representativeness, especially when the simulation output is used to inform policy or governance decisions [47].
Illusion of Consensus in Multi-Agent Systems.
Another downstream failure of social simulation is the illusion of consensus. Agents trained to average across pretraining data distributions may converge on responses that seem moderate or socially acceptable [48, 49], not because this reflects a reasoned position, but because it minimizes token-level loss [50, 51]. This leads to "simulated agreement" that aligns with a median perspective that masks underlying conflict, complexity, or division. In public policy contexts, this can produce misleading outcomes: agents might suggest a positive view of certain policies because the language model "agreed" in aggregate, when in fact no modeled agent went through the necessary deliberative reasoning to support that claim.
Motivating Example: In a simulated multi-agent townhall on climate adaptation strategies, a diverse group of generative agents representing rural residents, urban planners, coastal residents, and low-income renters are asked to deliberate on proposed flood insurance mandates. The agents converge on support for a single subsidy-based plan. Their justifications reference general ideas of “equity” and “resilience” as a safe middle-ground output but omit key concerns such as housing displacement, loss of land-based income, or municipal tax burdens.
As agents are increasingly used to test policy options, simulate deliberative processes, or represent groups in synthetic social systems, these reasoning deficiencies risk being institutionalized. Decision-makers may take model outputs at face value, unaware that these outputs do not derive from any concrete belief structure. Ultimately, when agents simulate without reasoning, the outputs they generate can erode trust, misinform policy, and flatten the epistemic diversity of the very populations they are meant to represent [44, 52, 53, 54].
3.2 Fidelity: Agents Should Be Coherent, Traceable, and Causally Grounded
Beyond demographic representation, simulations must ensure agents think in ways that are internally coherent, dynamically revisable, and causally structured. Current LLM-based agents fail to meet these standards. The mismatch between surface-level plausibility and internal structural fidelity manifests in a consistent set of reasoning failures:
Counterfactual Intervention Sensitivity and Belief Revision.
In social simulations, agents are expected to revise their stances when key assumptions or contextual conditions change–a hallmark of human reasoning known as counterfactual intervention sensitivity [55, 56]. However, current LLMs and LLM-based generative agents often respond to such interventions with inertia or token-level paraphrasing due to the lack of an internal causal structure that maps beliefs to causes and consequences [8, 57]. As a result, they cannot simulate the downstream effects of counterfactual changes, nor can they explain why a particular belief might hold under some conditions but not others.
This structural deficiency manifests as inconsistency across prompts or dialogue turns: an agent may support one policy in one scenario, then oppose in another, without any explicit causal revision or reasoning trace [58, 59]. While there have been research efforts in grounding models and model-based agents with causal memory [60] and knowledge graph [61], most of them focus on graph discovery and construction [62, 63] in specific knowledge domains rather than general human belief systems. Without an explicit model of how beliefs are formed, revised, and connected, agents’ utterances are generated in isolation—locally plausible, but globally incoherent.
Alternative View: Human cognition is non-monotonic and contextually fluid, thus demanding coherence is unrealistic.
One might argue that humans often hold incoherent or even contradictory beliefs. Demanding that agents simulate perfectly consistent beliefs risks idealizing cognition and misrepresenting actual human messiness [64, 65].
Response. We do not call for rigid logical coherence or monotonic reasoning. Rather, our claim is modest: agents should be able to faithfully simulate belief revision under counterfactual assumptions—not that they maintain perfect consistency across all scenarios. Furthermore, human belief systems can be messy and involve incoherent or even contradictory stances, but it doesn’t mean they’re structureless. Contradictions in actual human belief systems are often meaningful, reflecting a set of ambivalent conventions and priors in the social system at large [66]; in contrast, LLMs generate contradictions without memory, deliberation, or causal record. We don’t require logical perfection but rather grounded incoherence, that agents should be able to simulate how humans arrive at contradictory views and under what conditions those contradictions persist or resolve.
Traceability and Interpretability.
Beyond intervention sensitivity, current LLM-based agents also lack traceable and interpretable reasoning processes [67], which fundamentally limits their reliability and capacity for iterative improvement. These agents face faithfulness-related critiques similar to those observed in CoT-related discussions [68, 69]: they may produce fluent rationalizations after the fact, their "reasoning" is typically constructed post hoc—assembled from language patterns rather than derived from an underlying belief model [70, 71].
As a result, developers and researchers cannot inspect how a decision was reached, what assumptions it relied on, or how one belief led to another in a given context. Notably, there are some promising approaches beginning to incorporate structured representations, such as knowledge graphs, belief graphs, and additional reasoning layers, [72, 73, 74], but most of them are static, domain-specific, or disconnected from live generative processes; also, these emerging approaches haven’t been adopted by social simulations.
Alternative View: Post-hoc rationalization may be cognitively authentic and functionally sufficient.
Drawing from work in social and cognitive psychology, one might argue that people often rationalize decisions or beliefs after the fact [75, 76]. LLMs’ tendency to construct rationales retroactively might therefore be seen not as a defect, but as a cognitive parallel to human behaviors.
Response. We do not target post-hoc rationalization per se, but its total detachment from structured belief representation. Human justifications could be imperfect but still rely on internal models of causality, memory, and values and are traceable thereafter [77], whereas LLMs produce rationalizations without structured anchoring. Form (i.e. the shape of the rationale) is not the same as function (i.e. the structural, belief-guided deliberation). Agents must operate on traceable structures to simulate human reasoning.
Taken together, these are not isolated shortcomings but systemic indicators of structural misalignment. They reflect a foundational disconnect between the training objectives of large language models—next-token prediction over vast corpora—and the epistemic requirements of belief representation, causal inference, and deliberative revision [78]. Until this gap is addressed, generative agents will continue to exhibit the symptoms of alignment while remaining fundamentally unaligned at the level of thought.
3.3 Metric Illusion: Current Benchmarks are Far from Enough
Despite the growing sophistication of generative agents, evaluation standards remain stuck in a prior era of NLP—optimizing for stylistic fluency, short-term coherence, and plausibility of individual outputs. Most benchmarks treat language as a proxy for thought, assuming that coherent expression implies coherent reasoning. As a result, current evaluations risk rewarding agents for each response’s local plausibility, ignoring whether they maintain internal consistency across turns or scenarios. Stance classification tasks, for example, check whether a model picks a side but remain agnostic about how or why that stance was formed [79, 80]. Dialogue benchmarks reward conversational smoothness, even when agents flip positions over time [81, 82, 83]. These limitations flatten reasoning to performance—encouraging shallow persona simulation without testing whether agents have structured beliefs that guide their answers.
More critically, today’s benchmarks rarely assess belief revision or causal adaptability. Current evaluations assess agents on static inputs but neglect their responses to counterfactual interventions or other hypothetical changes. When prompted with counterfactuals (e.g., “What if surveillance were community-led?”), most models paraphrase prior stances rather than updating them through principled reasoning [84, 85, 86, 87, 88]. These failures reflect a deeper structural problem: we lack benchmarks that probe agents’ causal belief structures, trace their reasoning trajectories, or demand sensitivity to interventions.
4 What Does Human-Like Reasoning Entail?
The failures outlined above stem from a core mismatch between behavioral alignment and structural reasoning alignment. This section turns from diagnosis to design: what structural properties must agents possess to simulate human-like reasoning? We outline potential modeling paradigm shifts in social simulation, theoretical foundations drawn from cognitive science, and definitions of reasoning fidelity.
4.1 Modeling Paradigms in Social Simulation: A Cognitive Turn
While recent efforts in generative agent research focus on improving behavioral plausibility through techniques like persona prompting [21, 22], reinforcement learning from human feedback (RLHF) [31, 32], and chain-of-thought (CoT) generation [17, 89, 69], these methods share a common assumption: that plausible language implies plausible reasoning.
Our position challenges this assumption. These methods remain fundamentally output-centric, optimizing for stylistic fluency or stance alignment without simulating how beliefs are causally formed or revised. This often leads to post-hoc rationalizations, identity flattening, and the illusion of consensus.
By contrast, we propose a cognition-centric paradigm shift: modeling thought as a structured, revisable, and compositional process. Table 1 outlines this distinction.
Dimension Existing Paradigm Our Proposal (GenMinds)
Reasoning Format Token-level generation, post-hoc Structured belief graphs, motifs
Belief Dynamics Static or reset each prompt Revisable via causal updates
Evaluation Lens Output fluency, stance labels Reasoning fidelity and adaptability
Social Representation Averaged, flattened views Divergent, positional cognition
Table 1: Paradigm shift from output mimicry to cognitive modeling in generative agents.
4.2 Theoretical Foundations: Causal, Compositional, Revisable
To move beyond behavioral alignment, we must first define what it means to reason like a human.
Cognitive science offers a well-established answer. Decades of research suggest that human reasoning is not merely reactive output generation, but a process grounded in structured representations, counterfactual simulation, and dynamic belief updating [12, 14, 90]. From these foundations, we identify three defining features of human-like reasoning:
  1. 1.
    Causal: Humans reason in terms of causes and consequences. Even young children exhibit Bayesian-like inference over causal relationships and use interventions to test hypotheses about the world [12, 91]. Mental models are structured around “what caused what”—not just correlation, but explanation. This causal orientation allows for robust generalization and counterfactual reasoning [90].
  2. 2.
    Compositional: Human reasoning is modular and reusable. Cognitive architectures operate by composing shared schemas—what we term cognitive motifs—that generalize across domains [14, 16]. These motifs support efficient reasoning by enabling agents to simulate belief structures without re-learning from scratch [15].
  3. 3.
    Revisable: Human beliefs evolve dynamically. When presented with new information or contradiction, individuals revise their prior assumptions. This capacity for belief updating has been modeled through probabilistic programming and counterfactual simulation frameworks [15, 92], capturing the adaptive, non-monotonic nature of human thought.
Together, these three dimensions—causal, compositional, and revisable—form the foundation for what we call reasoning fidelity: the structural integrity of belief formation and revision processes in generative agents.
4.3 Defining Reasoning Fidelity
We define reasoning fidelity as an agent’s ability to construct, simulate, and revise a structured trace of belief formation that mirrors human causal reasoning patterns. This concept extends the dual-process model proposed by [92], in which language models interact with structured reasoning systems to model inference, belief, and decision-making.
Reasoning fidelity comprises three measurable properties:
  1. 1.
    Traceability — the ability to inspect how a belief or stance was formed through intermediate reasoning steps [93, 94];
  2. 2.
    Counterfactual adaptability — the capacity to revise beliefs predictably in response to interventions or changes in context [95, 96];
  3. 3.
    Motif compositionality — the reuse of modular causal structures (motifs) across different scenarios or domains [92, 97].
These properties define the core evaluation axes in the proposed RECAP paradigm, which shifts benchmarking from output plausibility to structural reasoning fidelity (Section 5). For example, traceability is assessed via motif-to-stance inference accuracy, adaptability through belief revision under hypothetical scenarios, and compositionality via motif reuse across unrelated topics.
This framework can be formed through explicit causal belief graphs, as illustrated in our proposed GenMinds architecture (Section 5). In such graphs, nodes represent causally relevant concepts (e.g., policy tradeoffs, values, or outcomes), and directed edges encode influence relationships. These graphs are derived from natural language using LLM-guided parsing and persist across interactions, enabling intervention analysis and reasoning trace reconstruction.
Importantly, this architecture is not tied to any particular implementation. While LLMs may serve as one plausible interface for extracting cognitive motifs, the core modeling contribution lies in structuring reasoning as revisable causal graphs—an approach compatible with both symbolic and neural systems [92]. GenMinds exemplifies one such instantiation of this broader modeling principle.
At the evaluation level, reasoning fidelity fulfills emerging demands for cognitively grounded AI benchmarks [84]. It offers a testable, interpretable standard for assessing agent behavior that goes beyond language mimicry.
Yet current LLM-based agents fall short of this standard. Most optimize for surface alignment—producing plausible stances like “I support policy X”—without modeling the underlying belief process. They lack persistent belief states, causal coherence, and principled revision under counterfactuals. This results in brittle or contradictory responses, agreement bias between agents, and an absence of traceable justification.
5 Toward Cognitively Grounded Simulation: Modeling and Evaluation Principles
After outlining the cognitive foundations necessary for human-like reasoning, we translate these principles into a modeling and evaluation framework for cognitively grounded simulation. This includes GenMinds, which models structured belief formation, and RECAP, which evaluates reasoning fidelity in generative agents.
5.1 GenMinds: A Framework for Modeling Human-Like Reasoning
Structured Thought Capture: From Semi-Structured Interviews to Causal Graphs.
To build generative agents that simulate human reasoning—not merely output plausible stances—we propose modeling individuals’ internal logic through semi-structured interviews, adaptively conducted by large language models (LLMs). These interviews elicit causal explanations in everyday language (e.g., “why do you support X?” “what does Y influence?”), which are then parsed into directed acyclic graphs representing the participant’s belief structure [3]. Each node encodes a concept (e.g., fairness, safety, family needs), and each edge encodes a directional causal relation, with confidence and polarity scores.
Shared Knowledge.
We introduce cognitive motifs as minimal causal reasoning units extracted from natural language. These motifs—e.g., “Surveillance → Crime Rate → Public Safety”—capture widely shared conceptual dependencies across individuals. When aggregated across interviews, they form a topology of commonly held belief structures.
We represent these motifs in a symbolic causal graph (CBN), enabling alignment of diverse opinions while maintaining transparency of reasoning. By grounding this structure in semi-structured interviews, we connect population-level reasoning to individual narratives.
Inference via Symbolic–Neural Hybrid Graph Simulation.
We define reasoning as a form of forward inference over belief graphs: given a causal structure and an intervention (e.g., “increasing housing near transit”), the agent uses probabilistic updates (e.g., do-calculus) to simulate belief shifts and final stances. A language model selects relevant interventions and assembles motifs into a causal Bayesian network. This hybrid method ensures both interpretability and expressive power, enabling agents to trace “why” a conclusion was reached and what would change it.
Be Aware of Unknown.
While causal motifs help model explicit reasoning patterns, real-world beliefs are often incomplete or contradictory. Our framework is designed to highlight missing links or uncertain dependencies by visualizing weakly supported or isolated nodes in the graph.
We encourage future systems to maintain uncertainty visualization and prompt-based elicitation to expand motif coverage, rather than overfitting to known paths. This allows belief modeling to remain adaptive and open-ended, rather than overly deterministic.
Figure 1: Motif-based belief graph and intervention. Natural language responses are parsed into motif-level causal links, forming a personalized belief graph. A simulated intervention on Transparency propagates downstream updates, shown as highlighted nodes and edges.
Illustrative Example: From Interviews to Reasoning Agent Structures
To concretize how motif-based causal reasoning operates in our framework, we present a real scenario from our semi-structured interviews on urban surveillance.
Step 1: Extracting causal motifs from QA responses.
We start with Q and A responses annotated with concept nodes and directional relations. For instance:
  • QA#1: Q: How do you think surveillance might affect public safety? A: “It can reduce crime by aiding investigations with more transparency, which increases public safety.” $\Rightarrow$ Motif: Transparency $\rightarrow$ Crime rate $\rightarrow$ Public safety
  • QA#2: Q: Does transparency have an effect on support for surveillance? A: “People will have less privacy concern if they know how data is used…” $\Rightarrow$ Motif: Privacy $\leftarrow$ Transparency $\rightarrow$ Crime rate
Step 2: Composing a Causal Belief Network.
These motifs are compiled into a belief graph representing the participant’s reasoning. Nodes are concepts; edges indicate directional influence. Confidence scores are derived from motif density or respondent emphasis.
Step 3: Simulating belief change via intervention.
We apply a hypothetical intervention:
do (Transparency = high)
This reflects a policy shift such as increasing camera accountability. Using belief propagation over the CBN, the downstream posteriors update as follows:
$P(\texttt{Privacy Concern})$ $:0.7\rightarrow 0.3$
$P(\texttt{Opposition to Surveillance})$ $:0.7\rightarrow 0.2$
This chain demonstrates the potential of motif-based causal modeling for simulating how real individuals might update beliefs when exposed to policy changes—moving beyond static opinion snapshots.
5.2 RECAP: Principles for Evaluating Reasoning Fidelity
To advance cognitively aligned simulation, we propose a benchmark framework—RECAP—that shifts evaluation from surface-level correctness to the internal structure and coherence of reasoning.
Design Principles.
  • Traceability: Can the agent construct a transparent chain of intermediate beliefs?
  • Demographic Sensitivity: Can it represent diverse reasoning paths across identities or contexts?
  • Intervention Coherence: Does it revise beliefs in response to hypothetical changes in a consistent, causally grounded way?
Structure and Inputs.
  • Situated prompt in a morally or socially complex domain;
  • Human-annotated responses capturing causal motifs and belief chains;
  • A task—e.g., graph reconstruction, stance explanation, or counterfactual reasoning—that requires structured inference.
Metrics.
  • Motif Alignment: Structural similarity between human and model belief graphs;
  • Belief Coherence: Internal consistency of the model’s reasoning trace;
  • Counterfactual Robustness: Sensible belief updates under interventions.
Grounding in Human Reasoning.
All items originate from real-world, semi-structured interviews, capturing how people explain and revise their beliefs. This grounding ensures the benchmark reflects the complexity and causal depth of actual human reasoning.
Toward a Shared Format.
RECAP is not a static dataset but a replicable schema for structured reasoning evaluation. Grounded in human-derived motifs, it aims to promote interpretability, adaptability, and socially responsible agent design.
6 A Call for Cognitively Grounded Simulation
As large language models become embedded in social simulations and policy tools, we face a pivotal choice: whether to pursue agents that merely sound human, or agents that can reason in structured, human-like ways. This paper argues for the latter. We call for a shift from behavior-level mimicry to cognitively grounded reasoning—where agents represent beliefs, simulate causal relationships, revise assumptions, and reveal their internal logic.
We introduced Generative Minds and RECAP as conceptual scaffolds to support this shift—prioritizing reasoning fidelity, traceability, and epistemic diversity over surface plausibility. These are not fixed systems, but a framework for developing agents that simulate how people think, not just what they say.
This paradigm enables more transparent diagnostics, pluralistic modeling of public reasoning, and structured evaluations that align with the complexity of real-world decisions.
Implications of Adopting Reasoning Fidelity as a Core Standard.
Adopting reasoning fidelity as a core standard would shift generative agent research from stylistic fluency to structural interpretability. It reshapes alignment evaluation, promotes modular and revisable architectures, and incentivizes cognitively grounded benchmarks. In high-stakes applications—civic simulation, participatory policy, AI governance—agents with causal transparency and revisable beliefs are essential for trust, auditability, and fairness. Without this shift, we risk institutionalizing brittle models that obscure bias and flatten the diversity of public thought.
What Comes Next.
We are actively developing:
  • Agent architectures for modular belief reasoning and counterfactual revision;
  • Tools for causal motif extraction and belief graph construction;
  • Datasets across domains such as housing, surveillance, and healthcare.
We invite the community to co-develop evaluation protocols, agent designs, and data pipelines that advance cognitively aligned simulation.
To simulate society faithfully, we must simulate thought.